OpenStudy (anonymous):

2) WILL GIVE MEDAL AND FAN!! Which of the following is not a perfect square trinomial? a. 169 – 26y + y2 b. 81 + 18y + y2 c. 64 + 8y + y2 d. 25 + 10y + y2

OpenStudy (anonymous):

which one do u think

OpenStudy (anonymous):

im not usre at all! @israa88

OpenStudy (anonymous):

i would go with A

OpenStudy (tkhunny):

ALL PSTs look like this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\) -- ALL of them. User the first and last term to predict the middle term

OpenStudy (anonymous):

so maybe A? @tkhunny

OpenStudy (tkhunny):

Maybe. Did you test it? See if it fits the form.

OpenStudy (anonymous):

ya I think it is A... @tkhunny

OpenStudy (anonymous):

you are right :)

OpenStudy (tkhunny):

You're guessing. Provide the square roots \(\sqrt{169} = \) \(\sqrt{81} = \) \(\sqrt{64} = \) \(\sqrt{25} = \)

OpenStudy (anonymous):

ok @tkhunny then what next ...

OpenStudy (tkhunny):

You didn't provide the square roots. What are they?

OpenStudy (tkhunny):

israa88 is also just guessing.

OpenStudy (anonymous):

lol @tkhunny honestly anyone can answer that question but that doesn't proof that they are guessing or not

OpenStudy (tkhunny):

I didn't say it did. I did say that you are guessing.

OpenStudy (tkhunny):

Now, where are those square roots?

OpenStudy (anonymous):

i am not guessing the answer to ur questions is 13, 9, 8 , and 5

OpenStudy (anonymous):

i am not guessing now am i

OpenStudy (tkhunny):

Sadly, now we don't know if brae could have done that. @brae You still have shown no work. Multiply those square roots by 2. What do you get?

OpenStudy (anonymous):

me or him ?

OpenStudy (tkhunny):

Who posted the question? @brae Are you still here? We're almost done.

OpenStudy (anonymous):

Well exactly Sir you can't come and say i am guessing as well because I really wasn't so have a wonderful day sir and next time plz dont judge :)

OpenStudy (anonymous):

umm this is kinda hard...

OpenStudy (tkhunny):

You're still guessing. Remember the form: \((a+b)^{2} = a^{2} + 2ab = b^{2}\) Constant terms: 169 81 64 25 -- These are the '\(b^{2}\)'s in the form. Square roots: 13 9 8 5 -- These are the '\(b\)'s in the form. Multiply by 2: 26 18 16 10 -- These represent "\(2b\)" Look at the coefficients on the middle terms. Did we hit them or not?